OpenStudy (anonymous):

please solve the equation by completing the square x^2+6x=1

OpenStudy (anonymous):

You need to complete the square to solve this problem. x^2-6x+1=0.....rewrite as x^2-6x=-1 square -6.....-6/2 = -3.....then -3^2 = 9.....add 9 both sides. x^2-6x+9 = -1+9........then, x^2-6x+9 = 8 then, factor right side.....(x-3)(x-3) which is (x-3)^2 then find the square root of 8......√8 which is ± 2.828.... Now find x: x-3 = +2.828.........add 3 to each side to isolate x. x-3+3 = +2.828+3....which is: x = 5.828 FIRST ANSWER. THEN, x-3 = -2.828.......add 3 to each side to isolate x. x-3+3 = -2.828+3....which is: x = 0.172 SECOND ANSWER. ANSWER IS: x = (0.172, 5.828) source: http://answers.yahoo.com/question/index?qid=20090424091700AADO6zV